# PositionIC: Unified Position and Identity Consistency for Image Customization

### v1

@@ Introduction
Some works (Wang et al. 2024b; Chen et al. 2024; Wang et al. 2024a; Jiménez 2023; Bar-Tal et al. 2023) have attempted to generate with precise layout control but few of them preserving the visual consistency where adhering layout instruction. Gligen (Li et al. 2023) encodes Fourier embedding as grounding token to inject position information. MS-diffusion (Wang et al. 2024c) utilizes grounding resampler correlating visual information with specific entities and spatial constraints. However, they still suffer from inconsistencies in vision and position, which hinders further application.
@@
Figure 7: Qualitative comparison of position control generation with different methods on PositionIC-Bench. MS-Diff and Gligen are existing position-controllable customization methods; RPF and Instance-Diff are position-only controllable methods.
@@
We compare our proposed PositionIC with several leading methods on DreamBench for both single-subject and multi-subject. As presented in Table 1 and Table 2, PositionIC achieves the highest scores on both CLIP-I and DINO of 0.846 and 0.828 on single-subject, 0.819 and 0.771 on multi-subject, respectively. PositionIC also has competitive CLIP-T compared to existing methods. The evaluation results indicate that PositionIC has remarkable performance on subject consistency and text fidelity.

### v2

@@ Introduction
Some works (Wang et al. 2024b; Chen et al. 2024; Wang et al. 2024a; Jiménez 2023; Bar-Tal et al. 2023; Shi et al. 2025) have attempted to generate with precise layout control. Gligen (Li et al. 2023) encodes Fourier embedding as grounding tokens to inject position information. MS-diffusion (Wang et al. 2024c) utilizes a grounding resampler correlating visual information with specific entities and spatial constraints. However, they still suffer from inconsistencies in vision and position, which hinder further application.
@@
Figure 7: Qualitative comparison of position control generation with different methods on PositionIC-Bench. RPF denotes Regional Prompting Flux. MS-Diff and Gligen are existing position-controllable customization methods; RPF and Instance-Diff are position-only controllable methods.
@@
We compare our proposed PositionIC with several leading methods on DreamBench for both single-subject and multi-subject. As presented in Table 1 and Table 2, PositionIC achieves the highest scores on both CLIP-I and DINO of 0.846 and 0.823 on single-subject, 0.819 and 0.771 on multi-subject, respectively. PositionIC also has competitive CLIP-T compared to existing methods. The evaluation results indicate that PositionIC has remarkable performance on subject consistency and text fidelity.